# Coriolis Forces Modify Magnetostatic Ponderomotive Potentials

### v1

@@ Abstract
It is possible to produce a ponderomotive effect in a plasma system without time-varying fields, if the plasma flows over spatially oscillations in the field. This can be achieved by superimposing a spatially oscillatory perturbation on a guide field, then setting up an electric field perpendicular to the guide field to drive flow over the perturbation. However, subtle distinctions in the structure of the resulting electric field can entirely change the behavior of the resulting ponderomotive force. Previous work has shown that, in slab models, these distinctions can be explained in terms of the polarization of the effective wave that appears in the co-moving frame. Here we consider what happens to this picture in a cylindrical system, where the transformation to the co-moving (rotating) frame is not inertial. It turns out that the non-inertial nature of this frame transformation can lead to counterintuitive behavior, partly due to the appearance of parallel (magnetic-field-aligned) electric fields in the rotating frame even in cases where none existed in the laboratory frame. Apart from the academic interest of this study, the practical impact lies in being better able to anticipate the antenna configuration on the plasma periphery of a cylindrical plasma that will lead to optimal ponderomotive barrier formation in the interior plasma.
@@ I Introduction
In the ponderomotive effect, a charged particle subject to an inhomogeneous wave field is repelled from or attracted to the region where the field is strongest. In the context of magnetic confinement for plasmas, there is longstanding interest in the possible use of ponderomotive forces to help trap plasmas. For plasma traps, the usual approach is to use an RF antenna to inject an electromagnetic wave into a particular region so as to better trap ions. [1, 2, 3, 4] This could be very helpful for improving confinement times for fusion devices. Much the same physics can be applied for applications involving plasma acceleration.[5] For some applications – for example, aneutronic fusion[6, 7, 8, 9, 10] – it may be of particular interest that these ponderomotive effects can selectively confine or deconfine different particle species.[11, 12, 13, 14]
@@
Figure 1: This cartoon illustrates the difference between two possible choices of electric field when the magnetic field is a straight guide field with a small perturbation. In the first case, the surfaces of constant voltage follow the unperturbed magnetic field lines. In the second, they follow the perturbed field lines.
@@ II The Role of Polarization
Figure 2: The qualitative structures of ponderomotive potentials assocaited with four different polarizations. The linear polarization is a combination of equal parts of the two circular polarizations. In all four cases, the amplitude of the electric field is assumed to be proportional to $\omega$.
@@
where $(x,y,z)$ are the usual Cartesian coordinates and the background magnetic field is in the $z$ direction. In a cold plasma, the resulting ponderomotive potential can be written as[21]
@@
where $q$ is the charge of the particle, $\omega$ is the frequency of the wave, and $\Omega\doteq qB/m$ is the gyrofrequency. Eq. (4) is only valid when the beat frequency varies slowly compared with the timescale of the field oscillations, so it breaks down near the singularities where $\Phi$ would otherwise diverge.[22] Formally, Eq. (4) can be used in a system with plasma flows by transforming to the co-moving (possibly non-inertial) frame. Fields with spatial oscillations in the laboratory frame can become time-varying in the co-moving frame.
@@ III Rotating-Frame Transformations
Formally, one way to calculate the behavior of the effecitve ponderomotive potential in a flowing plasma is to transform to the co-moving frame that eliminates the flow, then to calculate the decomposition of the electric field in that frame into $E_{+}$, $E_{-}$, and $E_{||}$. The resulting ponderomotive potential can be inferred from this decomposition in the same way described in the previous section. Throughout this paper, we will take drift velocities to be non-relativistic.
@@ IV.2 Higher-$n$ Multipole Fields
The square magnitude of $\mathbf{E}_{*}$ is
@@
so we can see that the smallness of the perturbation enters into $E_{*||}$ through $B_{1}/B_{0}$, whereas it enters into $E_{*}$ through $\xi=\int_{z_{0}}^{z}B_{1}\mathrm{d}z/RB_{0}$. As such, $E_{*||}\ll E_{*}$ whenever $B_{1}/B_{0}\ll\xi$ – or, equivalently, whenever the gradient scale length of $\xi$ is long compared with $R$. In these cases, the circularly-polarized component of $\mathbf{E}_{*}$ is much larger than the parallel-polarized component, and the ponderomotive pontential is well-approximated by the contribution from the circular-polarized component alone.
However, this does not mean that the parallel fields can be neglected. Although $E_{||}=0$ in the laboratory frame, and although a significant part of $E_{*||}$ is oscillatory, $E_{*||}$ does not average to zero over an oscillation. A field component that does not average to zero can contribute to the dynamics at a lower order than a comparably strong field component that does, because it can act on particles through direct acceleration (Joule heating) rather than through the ponderomotive effect.
So, when does this direct acceleration compete with the ponderomotive effect? Note that both terms in Eq. (42) contribute a nonzero term, on average. The first contributes though $\mathcal{O}(\xi)$ corrections to $r$ over the course of an oscillation; see Appendix A. The second, of course, is simply not oscillatory. The contribution from the nonoscillatory part of $E_{*||}$ comes from
$\Delta W_{\text{Joule}}\approx\int qE_{*||}\mathrm{d}z,$ (44)
since $\mathbf{B}_{*}$ is approximately parallel to $\hat{z}$. This amounts to the non-conservation of the form of the particle’s “quasi-energy” given in Eq. (1). Keeping in mind that $B_{1}/RB_{0}=\mathrm{d}\xi/\mathrm{d}z$, both of the two terms in Eq. (42) contribute $\mathcal{O}(\xi^{2})$ terms to $\Delta W_{\text{Joule}}$. This can be understood, at least in part, from the fact that the Coriolis force leads to a gap between flux surfaces and particle drift surfaces. As the magnetic field geometry guides a particle inwards or outwards, the Coriolis force systematically pushes the particle radially to regions with lower potential (electrostatic) energy. This is closely related to a mechanism of cross-field conductivity that appears in rotating systems.[24, 25]
@@
The key takeaways here are as follows. First, $E_{*||}$ is nonzero but appears at higher order in $\xi$ than the circularly-polarized component of $\mathbf{E}_{*}$. Therefore, the ponderomotive potential is dominantly set by the circularly-polarized wave that appears in the rotating frame rather than the parallel-polarized wave. Second, even though $E_{*||}$ is small, it does not average to zero. This means that (at leading order in $\xi$), $E_{*||}$ does not contribute to the ponderomotive potential $\Phi_{*}$, but it does contribute a direct acceleration term $\Delta W_{\text{Joule}}$ that appears at the same order as $\Phi_{*}$. Third, although $\Phi_{*}$ and $\Delta W_{\text{Joule}}$ are both $\mathcal{O}(\xi^{2})$, $\Delta W_{\text{Joule}}$ is suppressed relative to $\Phi_{*}$ in cases where $\omega_{\text{rot}}\ll\Omega_{0}$. This does not include all cases of interest, but it is the limit typically envisioned for fusion devices.
@@
Figure 3: Comparison of the effective potential barrier at $z=\ell$ between the low-frequency analytic predictions and numerics obtained using a particle-pusher. The analytics use the expression for the ponderomotive potential using the approximation that the field in the rotating frame is purely circularly polarized; see Eq. (4). As expected, we find that the circularly-polarized ponderomotive potential is a good approximation for the dynamics at low $\omega_{\text{rot}}/\Omega_{0}$ but a weaker approximation at higher $\omega_{\text{rot}}$.
@@ V Discussion
Our focus here has been exclusively on the question of how to understand the relationship between the fields and the resulting ponderomotive force (and specifically the role that inertial forces play in that relationship). There is, of course, another important problem that is closely related: how to determine what fields can actually be imposed on a real plasma system using practical coils or electrodes. This is a separate problem, and is discussed in more detail elsewhere.[14] In fact, it appears that the most promising practical candidate may be an X mode driven by electrostatic perturbations (as opposed to the magnetostatic perturbations envisioned in the example in Section IV).[14] These issues are related more generally to the problem of wave propagation in rotating plasmas. [29, 30, 31, 32, 33, 34, 35, 36, 37, 38] That being said, there is one comment on the wave-propagation problem that is worth making here. If we are in a limit where we expect $E_{||}$ to vanish – for example, due to the high mobility of charge carriers along field lines – it is reasonable to wonder whether it is the lab-frame $E_{||}$ or the rotating-frame $E_{*||}$ that ought to vanish. Note that the rotating-frame transformation depends on the charge and mass of the particle species being considered. The electrons are typically the most mobile charge carrier, and their low mass means that there is little difference between the parallel field in the different electron frames for them. In other words, if we are considering a limit where high charger-carrier mobility prevents parallel fields, it is sensible to take the laboratory-frame $E_{||}$ to vanish and to allow the ion rotating-frame $E_{*||}$ to be nonzero.
@@
All of the physics considered here is nonresonant, and indeed much of the analysis would not be valid too close to a resonance.[39] Static field perturbations can also be used to produce resonant wave-particle interactions; this was proposed by Fetterman and Fisch [40] in the context of $\alpha$ channeling.[41] The structure of the perpendicular-polarized ponderomotive force leads to stronger forces close to resonances, that is, as $-(n+2)\omega_{\text{rot}}$ or $(n-2)\omega_{\text{rot}}$ approach $\Omega_{0}$. For applications in which resonant effects are undesirable but strong ponderomotive forces are desirable, this leads to an interesting situation in which it is optimal to get as close to a resonance as possible without actually producing resonant interactions.

### v2

@@ Abstract
It is possible to produce a ponderomotive effect in a plasma system without time-varying fields, if the plasma flows over spatial oscillations in the field. This can be achieved by superimposing a spatially oscillatory perturbation on a guide field, then setting up an electric field perpendicular to the guide field to drive flow over the perturbation. However, subtle distinctions in the structure of the resulting electric field can entirely change the behavior of the resulting ponderomotive force. Previous work has shown that, in slab models, these distinctions can be explained in terms of the polarization of the effective wave that appears in the co-moving frame. Here we consider what happens to this picture in a cylindrical system, where the transformation to the co-moving (rotating) frame is not inertial. It turns out that the non-inertial nature of this frame transformation can lead to counterintuitive behavior, partly due to the appearance of parallel (magnetic-field-aligned) electric fields in the rotating frame even in cases where none existed in the laboratory frame. Apart from the academic interest of this study, the practical impact lies in being better able to anticipate the antenna configuration on the plasma periphery of a cylindrical plasma that will lead to optimal ponderomotive barrier formation in the interior plasma.
@@ I Introduction
In the ponderomotive effect, a charged particle subject to an inhomogeneous wave field is repelled from or attracted to the region where the field is strongest. In the context of magnetic confinement for plasmas, there is longstanding interest in the possible use of ponderomotive forces to help trap plasmas. For plasma traps, the usual approach is to use an RF antenna to inject an electromagnetic wave into a particular region so as to better trap ions. [1, 2, 3, 4] This could be very helpful for improving confinement times for fusion devices. Much the same physics can be applied for applications involving plasma acceleration.[5] For some applications – for example, aneutronic fusion[6, 7, 8, 9, 10] – it may be of particular interest that these ponderomotive effects can selectively confine or deconfine different particle species.[11, 12, 13, 14] Ponderomotive forces can also be used to stabilize otherwise unstable modes in magnetic confinement devices.[15, 16, 17, 18]
@@
Figure 1: This cartoon illustrates the difference between two possible choices of electric field when the magnetic field is a straight guide field with a small perturbation. In the first case, the surfaces of constant voltage (green) follow the unperturbed magnetic field lines (violet). In the second, they follow the perturbed field lines (so that the green and violet sit on top of one another).
@@ II The Role of Polarization
Figure 2: The qualitative structures of ponderomotive potentials associated with four different polarizations. The linear polarization is a combination of equal parts of the two circular polarizations. In all four cases, the amplitude of the electric field is assumed to be proportional to $\omega$. Note that, formally, either $\omega$ or $\Omega$ can be negative; $\Omega$ is the signed gyrofrequency and therefore depends on the sign of the particle’s charge, and the polarizations are defined in a way that allows the wave frequency to be negative (even though it might sometimes be simpler to map a negative-frequency wave with one circular polarization onto a positive-frequency wave with the other).
@@
where $(x,y,z)$ are the usual Cartesian coordinates and the background magnetic field is in the $z$ direction. In a cold plasma, the resulting ponderomotive potential can be written as[5, 15, 17, 18, 25]
@@
where $q$ is the charge of the particle, $\omega$ is the frequency of the wave, and $\Omega\doteq qB/m$ is the gyrofrequency. Eq. (4) breaks down near the singularities where $\Phi$ would otherwise diverge.[26] Formally, Eq. (4) can be used in a system with plasma flows by transforming to the co-moving (possibly non-inertial) frame. Fields with spatial oscillations in the laboratory frame can become time-varying in the co-moving frame.
@@ III Rotating-Frame Transformations
Formally, one way to calculate the behavior of the effective ponderomotive potential in a flowing plasma is to transform to the co-moving frame that eliminates the flow, then to calculate the decomposition of the electric field in that frame into $E_{+}$, $E_{-}$, and $E_{||}$. The resulting ponderomotive potential can be inferred from this decomposition in the same way described in the previous section. Throughout this paper, we will take drift velocities to be non-relativistic.
@@ IV.2 Higher-$n$ Multipole Fields
The sign of the first term in Eq. (42) depends on the correlation of $r$ with $\theta$; see Appendix A. The second term is negative if $E_{0}$, $B_{1}$, and $\xi$ are positive. The square magnitude of $\mathbf{E}_{*}$ is
@@
so we can see that the smallness of the perturbation enters into $E_{*||}$ through $B_{1}/B_{0}$, whereas it enters into $E_{*}$ through $\xi=\int_{z_{0}}^{z}B_{1}\mathrm{d}z/RB_{0}$. As such, $E_{*||}\ll E_{*}$ whenever $B_{1}/B_{0}\ll\xi$ – or, equivalently, whenever the gradient scale length of $\xi$ is long compared with $R$. In these cases, the circularly-polarized component of $\mathbf{E}_{*}$ is much larger than the parallel-polarized component, and the ponderomotive potential is well-approximated by the contribution from the circular-polarized component alone.
However, this does not mean that the parallel fields can be neglected. Although $E_{||}=0$ in the laboratory frame, and although a significant part of $E_{*||}$ is oscillatory, $E_{*||}$ does not average to zero over an oscillation. A field component that does not average to zero can contribute to the dynamics at a lower order than a comparably strong field component that does, because it can act on particles through direct acceleration rather than through the ponderomotive effect.
So, when does this direct acceleration compete with the ponderomotive effect? Note that both terms in Eq. (42) contribute a nonzero term, on average. The first contributes $\mathcal{O}(\xi)$ corrections to $r$ over the course of an oscillation; see Appendix A. The second, of course, is simply not oscillatory. The contribution from the nonoscillatory part of $E_{*||}$ comes from
$\Delta W_{\text{direct}}\approx\int qE_{*||}\mathrm{d}z,$ (44)
since $\mathbf{B}_{*}$ is approximately parallel to $\hat{z}$. This amounts to the non-conservation of the form of the particle’s “quasi-energy” given in Eq. (1). Keeping in mind that $B_{1}/RB_{0}=\mathrm{d}\xi/\mathrm{d}z$, both of the two terms in Eq. (42) contribute $\mathcal{O}(\xi^{2})$ terms to $\Delta W_{\text{direct}}$. This can be understood, at least in part, from the fact that the Coriolis force leads to a gap between flux surfaces and particle drift surfaces. As the magnetic field geometry guides a particle inwards or outwards, the Coriolis force systematically pushes the particle radially to regions with lower potential (electrostatic) energy. This is closely related to a mechanism of cross-field conductivity that appears in rotating systems.[28, 29]
@@
The key takeaways here are as follows. First, $E_{*||}$ is nonzero but appears at higher order in $\xi$ than the circularly-polarized component of $\mathbf{E}_{*}$. Therefore, the ponderomotive potential is dominantly set by the circularly-polarized wave that appears in the rotating frame rather than the parallel-polarized wave. Second, even though $E_{*||}$ is small, it does not average to zero. This means that (at leading order in $\xi$), $E_{*||}$ does not contribute to the ponderomotive potential $\Phi_{*}$, but it does contribute a direct acceleration term $\Delta W_{\text{direct}}$ that appears at the same order as $\Phi_{*}$. Third, although $\Phi_{*}$ and $\Delta W_{\text{direct}}$ are both $\mathcal{O}(\xi^{2})$, $\Delta W_{\text{direct}}$ is suppressed relative to $\Phi_{*}$ in cases where $\omega_{\text{rot}}\ll\Omega_{0}$. This does not include all cases of interest, but it is the limit typically envisioned for fusion devices.
@@
Figure 3: Comparison of the effective potential barrier at $z=\ell$ between the low-frequency analytic predictions and numerics obtained using a particle-pusher. The analytics use the expression for the ponderomotive potential using the approximation that the field in the rotating frame is purely circularly polarized; see Eq. (4). As expected, we find that the circularly-polarized ponderomotive potential is a good approximation for the dynamics at low $\omega_{\text{rot}}/\Omega_{0}$ but a weaker approximation at higher $\omega_{\text{rot}}$. The shaded backgrounds provide a reference scale for the validity of the approximation, with darker gray appearing when $\omega_{\text{rot}}$ becomes less small compared with $\Omega_{0}$.
@@
In general, $E_{*||}$ may take either sign; depending on the parameters of the system, the parallel acceleration can either increase or reduce the effects of the ponderomotive potential. In the examples shown in Figure 3, its most noticeable effect seems to be to reinforce the ponderomotive potential when $(n-2)\omega_{\text{rot}}>\Omega_{0}$ (to the right of the resonance).
@@ V Discussion
Our focus here has been exclusively on the question of how to understand the relationship between the fields and the resulting ponderomotive force (and specifically the role that inertial forces play in that relationship). There is, of course, another important problem that is closely related: how to determine what fields can actually be imposed on a real plasma system using practical coils or electrodes. This is a separate problem, and is discussed in more detail elsewhere.[14] In fact, it appears that the most promising practical candidate may be an X mode driven by electrostatic perturbations (as opposed to the magnetostatic perturbations envisioned in the example in Section IV).[14] These issues are related more generally to the problem of wave propagation in rotating plasmas. [33, 34, 35, 36, 37, 38, 39, 40, 41, 42] That being said, there is one comment on the wave-propagation problem that is worth making here. If we are in a limit where we expect $E_{||}$ to vanish – for example, due to the high mobility of charge carriers along field lines – it is reasonable to wonder whether it is the lab-frame $E_{||}$ or the rotating-frame $E_{*||}$ that ought to vanish. Note that the rotating-frame transformation depends on the charge and mass of the particle species being considered. The electrons are typically the most mobile charge carrier, and their low mass means that there is little difference between the parallel field in the different electron frames for them. In other words, if we are considering a limit where high charge-carrier mobility prevents parallel fields, it is sensible to take the laboratory-frame $E_{||}$ to vanish and to allow the ion rotating-frame $E_{*||}$ to be nonzero.
@@
All of the physics considered here is nonresonant, and indeed much of the analysis would not be valid too close to a resonance.[43] Static field perturbations can also be used to produce resonant wave-particle interactions; this was proposed by Fetterman and Fisch [44] in the context of $\alpha$ channeling.[45] The structure of the perpendicular-polarized ponderomotive force leads to stronger forces close to resonances, that is, as $-(n+2)\omega_{\text{rot}}$ or $(n-2)\omega_{\text{rot}}$ approach $\Omega_{0}$. Correctly simulating the behavior very near resonance can be a subtle problem, and is likely to benefit from advanced computational techniques.[46] For applications in which resonant effects are undesirable but strong ponderomotive forces are desirable, this leads to an interesting situation in which it is optimal to get as close to a resonance as possible without actually producing resonant interactions.
Also note that a helical magnetic perturbation could produce an effective wave with a nonzero wavenumber component in the parallel direction. One could consider, for example, the helical structure imposed on the SMOLA device at the Budker Institute.[47] Such a perturbation would duly modify the potential described in Eq. (14); the resulting Doppler shift would depend on $v_{||}$, the parallel component of the particle’s velocity.
Of course, when considering a species-dependent effect like this one, it is important to keep in mind that the overall response of a plasma system will include both the effect itself and the self-consistent response of the plasma to the effect. A helpful analogy might be the trapping effect in a magnetic mirror: magnetic mirrors more effectively trap ions than electrons, so the imposition of a mirror field leads to an ambipolar electric field that tends to equalize the overall confinement times for ions and electrons. When considering the behavior of the plasma as a whole, it is sometimes practical to work with a combined ponderomotive force, summed over multiple species.[17, 18] This would be reasonably straightforward to do for the case considered here, though it would be important to keep in mind that parameters like the rotation frequency are also species-dependent.